What is the present value of $10,000 received a, 12 years from today when the interest rate is 4% per year? b, 20 years from today when the interest rate is 8% per year? 4. c, 6 years from today when the interest rate is 2% C.
Answer #1

Present Value = cash flow at period/(1+rate of return)^number of periods

A. PV= 10,000/(1.04)^12= $6245.97

B. PV = 10,000/(1.08)^20=$2145.48

C. PV = 10,000/(1.02)^6=$8879.71